OpenStudy (cutiecomittee123):

Solve this system of conic section equations 7y^2+x^2=64 x+y=4

jimthompson5910 (jim_thompson5910):

First solve the equation x+y = 4 for y \[\Large x+y=4\] \[\Large y=4-x\] Now plug this into the other equation so you boil things down to one variable only \[\Large 7y^2+x^2=64\] \[\Large 7(4-x)^2+x^2=64\] \[\Large 7(16-8x+x^2)+x^2=64\] \[\Large 112-56x+7x^2+x^2=64\] \[\Large 112-56x+8x^2=64\] \[\Large 112-56x+8x^2-64=0\] \[\Large 8x^2-56x+48=0\] From this point, you need to solve for x. To do so, I recommend using the quadratic formula. I'll let you do this part. Tell me what x values you get as the solutions.

OpenStudy (cutiecomittee123):

I get x=1 and x=6

jimthompson5910 (jim_thompson5910):

me too

jimthompson5910 (jim_thompson5910):

if x = 1, then y = ???

OpenStudy (cutiecomittee123):

sweet and then just plug that into on of the original equations to solve for y

jimthompson5910 (jim_thompson5910):

yes or plug into y = 4-x since that already has y isolated

OpenStudy (cutiecomittee123):

so like y=4-1 and y=4-6

OpenStudy (cutiecomittee123):

y=3 and y=-2

jimthompson5910 (jim_thompson5910):

correct, so the solutions are these two ordered pairs (1,3) and (6,-2) and this confirms it http://www.wolframalpha.com/input/?i=7y^2%2Bx^2%3D64%2Cx%2By%3D4

OpenStudy (cutiecomittee123):

wanna help me with another one?

jimthompson5910 (jim_thompson5910):

sure

jimthompson5910 (jim_thompson5910):

what's your question?

OpenStudy (cutiecomittee123):

x^2+y^2+2x+2y=0 x^2+y^2+4x+6y+12=0 solve this system of equations

jimthompson5910 (jim_thompson5910):

ok at first this looks really complicated, but we can eliminate quite a bit here notice how each equation has x^2+y^2 in it so we can subtract the equations (either equation1 - equation2 or equation2-equation1) to eliminate the x^2+y^2 terms what do you get when you subtract?

OpenStudy (cutiecomittee123):

2x+4y+12=0

jimthompson5910 (jim_thompson5910):

Now let's solve 2x+4y+12=0 for x 2x+4y+12=0 2x+4y+12-12=0-12 2x+4y = -12 2x+4y-4y = -12 - 4y 2x = -4y - 12 2x/2 = (-4y-12)/2 x = -2y - 6

OpenStudy (cutiecomittee123):

now plug that into the other equation

jimthompson5910 (jim_thompson5910):

Next, plug x = -2y - 6 into either original equation. I'll pick the first equation x^2+y^2+2x+2y=0 (-2y-6)^2+y^2+2(-2y-6)+2y=0 ... replace x with -2y-6 (4y^2+24y+36)+y^2+2(-2y-6)+2y=0 4y^2+24y+36+y^2-4y+12+2y=0 5y^2+22y+48=0 solve that for y (use the quadratic formula). Tell me what you get

jimthompson5910 (jim_thompson5910):

yes correct

jimthompson5910 (jim_thompson5910):

hmm I messed up, let me fix

jimthompson5910 (jim_thompson5910):

x^2+y^2+2x+2y=0 (-2y-6)^2+y^2+2(-2y-6)+2y=0 ... replace x with -2y-6 (4y^2+24y+36)+y^2+2(-2y-6)+2y=0 4y^2+24y+36+y^2-4y-12+2y=0 ... it should be -12, not +12 5y^2+22y+24=0

OpenStudy (cutiecomittee123):

well I tried and I got the b^2-4(a)(c) = -476 you cant take a square root of a negative number

jimthompson5910 (jim_thompson5910):

I fixed my mistake and got 5y^2+22y+24=0

OpenStudy (cutiecomittee123):

oh gotcha let me try that

OpenStudy (cutiecomittee123):

y=-2 and y=-2.4

jimthompson5910 (jim_thompson5910):

-2.4 or -12/5

jimthompson5910 (jim_thompson5910):

now use each y value to find the corresponding x value

OpenStudy (cutiecomittee123):

-12/5? how did you get that?

OpenStudy (cutiecomittee123):

-22+ or =- the sqrt of 4

OpenStudy (cutiecomittee123):

over 10

jimthompson5910 (jim_thompson5910):

I used the quadratic formula or you can convert -2.4 to fraction form -2.4 = -2.4*(10/10) = -24/10 = -12/5

OpenStudy (cutiecomittee123):

makes sense

OpenStudy (cutiecomittee123):

so now we plug in the y values to get x

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

you can use x = -2y - 6

OpenStudy (cutiecomittee123):

x=-10.8 and x=-10

jimthompson5910 (jim_thompson5910):

incorrect on both

jimthompson5910 (jim_thompson5910):

if y = -2, then x = -2y - 6 x = -2(-2) - 6 x = 4 - 6 x = -2

jimthompson5910 (jim_thompson5910):

if y = -2.4, then x = -2y - 6 x = -2(-2.4) - 6 x = 4.8 - 6 x = -1.2

OpenStudy (cutiecomittee123):

so the solutions are (-2,-2) and (-1.2, -12/

jimthompson5910 (jim_thompson5910):

correct and the graph confirms it

jimthompson5910 (jim_thompson5910):

I think you meant to say -12/5

jimthompson5910 (jim_thompson5910):

if you use -12/5 then convert -1.2 to fraction form as well -1.2 = -1.2*(10/10) = -12/10 = -6/5

OpenStudy (cutiecomittee123):

Yeah I did mean -12/5 thanks:))

jimthompson5910 (jim_thompson5910):

np